# Boosting Semi-Supervised 2D Human Pose Estimation by Revisiting Data Augmentation and Consistency Training

### v2

@@ 3.1 Paradigms of Generating Superior Augmentations
Ranking of Basic Augmentations. The core of the easy-hard pair paradigm $(\mathbf{I}_{e},\mathbf{I}_{h})$ is a more advanced augmentation. For this reason, Dual-Network [58] and SSPCM [23] propose pseudo keypoint-based augmentations Joint Cutout ($T_{JC}$) and Joint Cut-Occlude ($T_{JO}$), respectively. They also reach a similar yet crude conclusion about difficulty levels of existing augmentations: $\{T_{JO},T_{JC}\}\!>\!\{T_{RA},T_{CM},T_{CO},T_{MU},T_{A60}\}$, where $T_{RA}$, $T_{CM}$, $T_{CO}$ and $T_{MU}$ are RandAugment [11], CutMix [66], Cutout [12] and Mixup [70], respectively. The $T_{A60}$ consists of two $T_{A30}$. We give them a new ranking by conducting more rigorous trainings one-by-one. The $T_{RA}$ is removed for it contains repetitions with $T_{CO}$ and $T_{A60}$. As shown in Fig. 3, we divide the rest by their distinguishable gaps into four levels: $\{T_{JC},T_{JO}\}\!>\!\{T_{CO},T_{CM}\}\!>\!\{T_{MU}\}\!>\!\{T_{A60}\}$.
@@
Synergy between Augmentations. Then, instead of laboriously designing stronger augmentations, we consider to conduct two or more augmentations in sequence to obtain superior combinations conveniently. This idea is feasible because it essentially belongs to the AutoAug families [10, 33, 18, 72]. Instead of auto-searching, we expect to find some heuristic strategies for the HPE task. In fact, after performing joint-related $T_{JO}$ or $T_{JC}$ on one image, we can continue to perform some joint-unrelated augmentations such as $T_{CM}$, $T_{CO}$ and $T_{MU}$ on random areas. As shown in Fig. 3, applying $T_{JOCO}$ (a $T_{CO}$ after $T_{JO}$) or $T_{JCCM}$ (a $T_{CM}$ after $T_{JC}$) will bring harder samples for generating more effective student signals, but not destroy the semantic information visually. We call this discovery the synergistic effect between different augmentations. The $T_{A60}$ can server as an essential factor in any $T_{h}$ for keeping the geometric diversity.
@@
Selection of Augmentations Combination. Now, if selecting from the rest five basic augmentations, there are up to 26 choices $(2^{5}-\binom{5}{0}-\binom{5}{1})$. The optimal combinations are still time-consuming to acquire. Fortunately, not arbitrary number or kind of augmentations are collaborative. We intuitively summarize three simplistic principles. (P1) A global $T_{MU}$ does not make sense for the HPE task. (P2) Stacking augmentations with the similar perturbation type (e.g., $T_{JO}\!\sim\!T_{CM}$ and $T_{JC}\!\sim\!T_{CO}$) or difficulty level (e.g., $T_{JO}\!\sim\!T_{JC}$ and $T_{CM}\!\sim\!T_{CO}$) may not bring significant gain. (P3) Adding too many augmentations (e.g., three or four) will be profitless or even harmful for seriously polluting the image. We thus nominate the most likely superior combinations: $T_{JOCO}$ and $T_{JCCM}$. A case of setting $T_{e}$ as $T_{A30}$ and $T_{h}$ as $T_{JOCO}$ for getting corresponding easy teacher signals $\mathbf{H}_{e}$ and hard student signals $\mathbf{H}_{h}$ is shown as below:
$\mathbf{H}_{e}=T_{A30\rightarrow A60}(f(\mathbf{I}_{e})),\quad\mathbf{I}_{e}=T_{A30}(\mathbf{I})$ (3)
$\mathbf{H}_{h}=f(\mathbf{I}_{h}),\quad\mathbf{I}_{h}=T_{A30}(T_{CO}(T_{JO}(T_{A30}(\mathbf{I}))))~{}$
@@
where $T_{A60}$ is default to $T_{h}$, and divided into two separate $T_{A30}$ for being compatible with $T_{e}$. To further verify the above intuitive principles, we follow the augmentations ranking way and conduct empirical studies on the performance of up to 13 selected representative combinations out of 26 choices.
Table 1: Best mAP results of different combinations tests.
@@
Figure 4: The convergence curves of different tests in Tab. 3.1.
Table 2: Best mAP results of different MultiAugs tests.
@@
$m_{9}$ $T_{JCCM}$, $T_{JCCM}$ 43.6
$m_{10}$ $T_{JOCO}$, $T_{JOCO}$ 44.0
@@
Figure 5: The convergence curves of different tests in Tab. 3.1.
As shown in Tab. 3.1 and Fig. 3.1, we can examine three principles one-by-one: (P1) $T_{MU}$ often causes adverse or inferior effects for each combination. Please refer paired combinations $c_{1}$-$c_{3}$, $c_{2}$-$c_{4}$, $c_{1}$-$c_{8}$ and $c_{2}$-$c_{9}$. Thus, we do not add it for finding superior combinations. (P2) Synergistic effects between augmentations do exist. Please refer paired combinations $c_{1}$-$c_{5}$, $c_{1}$-$c_{7}$, $c_{2}$-$c_{6}$ and $c_{2}$-$c_{7}$. Especially, the $T_{JCJO}$ with two most advanced augmentations performs the worst among combinations {$c_{1}$,$c_{2}$,$c_{5}$,$c_{6}$,$c_{7}$}, which roundly reveals the harm of violating the principle of synergy. (P3) Do not overly combine too many augmentations. Please refer paired combinations $c_{1}$-$c_{12}$, $c_{2}$-$c_{13}$, $c_{1}$-$c_{11}$ and $c_{2}$-$c_{10}$. Stacking more augmentations brings non-significant gains or even results in degradation. We attribute it to deviating from the rule of collaboration and possibly producing meaningless or difficult-to-recognize images. Based on these facts, we have sufficient reasons not to check the performance of rest combinations. And we recommend two new strongest combinations $T_{JOCO}$ and $T_{JCCM}$.
3.2 Multi-path Consistency Losses
$\small\mathcal{L}^{*}_{u}=\mathbb{E}_{\mathbf{H}_{h_{i}}\in\mathcal{H}^{n}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e}-\mathbf{H}_{h_{i}}||^{2}~{}$ (4)
where $\mathbf{H}_{e}$ and $\mathbf{H}_{h_{i}}$ are obtained as dissected in Eq. 3. The $\mathbf{H}_{e}$ keeps constant for each $\mathbf{H}_{h_{i}}$. Comparably, Data Distill [45] applies a single model to multiple transformations of unlabeled data to train a student model. Then, it ensembles predictions to obtain keypoint locations, and re-generates a pseudo heatmap for supervision. Differently, we argue that conducting a fusion on predicted heatmaps in SSHPE is harmful. We consider that there are always differences in the estimation of keypoint positions for each $\mathbf{I}_{h_{i}}$. It is an ill-posed problem to heuristically evaluate the consistency regularization contribution of each heatmap in pixel during ensemble. We will explain this in ablation studies Sec. 5.4.
Despite the simplicity, such a minor modification brings consistent gains over the original Single-Network under same SSL settings. With our discovered augmentation combinations $T_{JOCO}$ and $T_{JCCM}$, the boosted Single-Network can surpass the original Dual-Network evidently. We validated in ablation studies that the performance gain is non-trivial. We conjecture that regularizing multiple hard augmentations with a shared easy augmentation can be regarded as enforcing consistency among advanced augmentations as well, which inherits the concept of training positive-negative paired samples in contrastive learning [6, 19, 7] and its SSL-related variations [30, 62, 54].
Effectiveness Verification. As shown in Tab. 3.1 and Fig. 3.1, we also experimentally verified the major advantage of the multi-path augmentations (dubbed as MultiAugs) strategy. Here, we have two variables of MultiAugs: the number of paths and the category of augmentations. For acquisition of an optimal augmentations set, similarly, we continue with the rules summarized in the previous section, and elect 12 different MultiAugs schemes for illustrating. We can witness the effectiveness of MultiAugs from two aspects: (1) It can inherit and even expand the synergistic effects between different augmentations. (2) It can alleviate the defects caused by excessive stacking of augmentations.
@@ 3.2 Multi-path Consistency Losses
Specifically, by comparing $c_{1}$-$m_{1}$, $c_{2}$-$m_{2}$ and $c_{7}$-$m_{2}$, we find that MultiAugs provides comparable performance with the same augmentations. By comparing $c_{10}$-$m_{6}$, $c_{11}$-$m_{7}$, $c_{12}$-$m_{4}$ and $c_{13}$-$m_{5}$, we can observe the sustained large gains brought by MultiAugs. The scheme $m_{8}$ does not obtain a better result than $m_{11}$ showing that new augmentations $T_{JOCO}$ and $T_{JCCM}$ have their essential properties. Schemes $m_{9}$ and $m_{10}$ which utilize a single augmentation twice are also inferior to $m_{11}$ showing the cooperativity of using multi-path distinct augmentations. Moreover, the optimal scheme $m_{12}$ further unleashes capabilities of four advanced augmentations. In order to balance performance and time consumption, we do not add more augmentation paths.
3.3 Analysis of Superior Augmentations
(a) Top-50 singular values
(b) mAP vs. Entropy
Figure 6: SVD analysis. The backbone of all six models is ResNet18. The 512-D features of 6,352 samples in COCO val-set are extracted.
@@ 3.3 Analysis of Superior Augmentations
In this part, we tentatively analyze why employing a superior augmentation to strongly augment the unlabeled data can improve model performance. Different from UDA [56] using the improved connectivity of constructed graphs to explain, we start from the perspective of shaped feature space of unlabeled data based on the singular value spectrum, which is widely considered to be related to the model transferability and generalization [8, 60]. Specifically, we perform singular value decomposition (SVD) on features $\mathbf{F}\!\in\!\mathbb{R}^{N\times D}$ [^1] We denote $N$ as the number of samples and $D$ as feature dimensions (a.k.a, $D\leq N$). extracted by various trained models with different strong augmentations on one dataset: $\mathbf{F}\!=\!\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, where $\mathbf{U}$ and $\mathbf{V}$ is the left and right singular vector matrices, respectively, and $\mathbf{\Sigma}$ denotes the diagonal singular value matrix $\{\sigma_{1},\sigma_{2},...,\sigma_{D}\}$. Then, we plot calculated singular values in Fig. 5(a). To further measure the flatness of the singular value distribution, we calculate the entropy of normalized singular values $\mathsf{H}_{nsv}$:
$\small\mathsf{H}_{nsv}=-\sum\nolimits^{D}_{i=1}\frac{\sigma_{i}}{\sum\nolimits^{D}_{j=1}\sigma_{j}}\log{\frac{\sigma_{i}}{\sum\nolimits^{D}_{j=1}\sigma_{j}}}~{}$ (5)
@@
Usually, a larger $\mathsf{H}_{nsv}$ indicates that the feature space captures more structure in the data and thus spans more dimensions due to more discriminated representations learned. As shown in Fig. 5(b), the model performance is positively correlated with the $\mathsf{H}_{nsv}$ value. Therefore, a superior augmentation facilitates better model generalization to unseen test sets.
4 Our Overall Framework: MultiAugs
We leverage unlabeled images by applying multiple augmentations with integrating two key techniques introduced in Sec. 3.1 and Sec. 3.2. Firstly, assuming that we have obtained an optimal augmentation set $\mathcal{\widehat{T}}^{n}\!=\!\{\widehat{T}_{h_{i}}|^{n}_{i=1}\}$, where $\widehat{T}_{h_{i}}$ may be an old single augmentation or a novel combined one. Then, we present how to construct our overall training framework in Fig. 7 based on either the Single-Network or the Dual-Network.
(a) MultiAugs - Single Network
(b) MultiAugs - Dual Networks
@@ 4 Our Overall Framework: MultiAugs
Figure 7: Our MultiAugs uses the Single Network (a) or Dual Networks (b), which can utilize multiple hard augmentations and facilitate multi-path consistency training.
@@
MultiAugs (Dual-Network). As shown in Fig. 6(b), this framework learns two identical yet independent networks with each similar to the Single-Network. For one input batch in every step, each of the two networks serves as both a teacher and a student. They both are fed by easy and hard augmentations of unlabeled images $\mathbf{I}^{u}$ when they produce teacher signals and student signals, respectively. Assuming we have two networks $f_{A}$ and $f_{B}$, and also the augmented easy images $\mathbf{I}^{u}_{e}$ using $T_{e}$ and $n$-path hard images $\{\mathbf{I}^{u}_{h_{1}},\mathbf{I}^{u}_{h_{2}},...,\mathbf{I}^{u}_{h_{n}}\}$ using $\mathcal{\widehat{T}}^{n}$, we first predict the following four types of heatmaps:
$\mathbf{H}_{e,A}=T_{A30\rightarrow A60}(f_{A}(\mathbf{I}^{u}_{e})),\quad\mathcal{H}_{A}=\{\mathbf{H}_{h_{i},A}|^{n}_{i=1},\mathbf{H}_{h_{i},A}=f_{A}(\mathbf{I}^{u}_{h_{i}})\}$ (6)
$\mathbf{H}_{e,B}=T_{A30\rightarrow A60}(f_{B}(\mathbf{I}^{u}_{e})),\quad\mathcal{H}_{B}=\{\mathbf{H}_{h_{i},B}|^{n}_{i=1},\mathbf{H}_{h_{i},B}=f_{B}(\mathbf{I}^{u}_{h_{i}})\}$

### v3

@@ III-A1 Ranking of Basic Augmentations
The core of the easy-hard pair paradigm $(\mathbf{I}_{e},\mathbf{I}_{h})$ is an advanced augmentation. To this end, Dual-Network [17] and SSPCM [20] propose pseudo keypoint-based augmentations Joint Cutout ($T_{JC}$) and Joint Cut-Occlude ($T_{JO}$), respectively. They also reach a similar yet crude conclusion about difficulty levels of existing augmentations: $\{T_{JO},T_{JC}\}\!>\!\{T_{RA},T_{CM},T_{CO},T_{MU},T_{A60}\}$, where $T_{RA}$, $T_{CM}$, $T_{CO}$ and $T_{MU}$ are RandAugment [26], CutMix [24], Cutout [22] and Mixup [23], respectively. The $T_{A60}$ consists of two $T_{A30}$. We give them a new ranking by conducting more rigorous trainings one-by-one. The $T_{RA}$ is removed for it contains repetitions with $T_{CO}$ and $T_{A60}$. As shown in Fig. 2, we divide the rest by their distinguishable gaps into four levels: $\{T_{JC},T_{JO}\}\!>\!\{T_{CO},T_{CM}\}\!>\!\{T_{MU}\}\!>\!\{T_{A60}\}$.
Figure 3: Illustrations of superior combinations $T_{JOCO}$ and $T_{JCCM}$. Either of them is a sequential operations of ready-made collaborative augmentations. $T_{JO}$ and $T_{CM}$ need extra patches cropped from other images which are not displayed.
III-A2 Synergy between Augmentations
@@
Then, instead of laboriously designing stronger augmentations, we consider to conduct two or more augmentations in sequence to obtain superior combinations conveniently. This idea is feasible because it essentially belongs to the AutoAug families [25, 27, 28, 29]. Instead of auto-searching, we expect to find some heuristic strategies for the HPE task. In fact, after performing joint-related $T_{JO}$ or $T_{JC}$ on one image, we can continue to perform some joint-unrelated augmentations such as $T_{CM}$, $T_{CO}$ and $T_{MU}$ on random areas. As shown in Fig. 3, applying $T_{JOCO}$ (a $T_{CO}$ after $T_{JO}$) or $T_{JCCM}$ (a $T_{CM}$ after $T_{JC}$) will bring harder samples for generating more effective student signals, but not destroy the semantic information visually. We call this discovery the synergistic effect between different augmentations. The $T_{A60}$ can server as an essential factor in any $T_{h}$ for keeping the geometric diversity.
III-A3 Selection of Augmentations Combination
@@
Now, if selecting from the rest five basic augmentations, there are up to 26 choices $(2^{5}-\binom{5}{0}-\binom{5}{1})$. The optimal combinations are still time-consuming to acquire. Fortunately, not arbitrary number or kind of augmentations are collaborative. We intuitively summarize three simplistic principles. (P1) A global $T_{MU}$ does not make sense for the HPE task. (P2) Stacking augmentations with the similar perturbation type (e.g., $T_{JO}\!\sim\!T_{CM}$ and $T_{JC}\!\sim\!T_{CO}$) or difficulty level (e.g., $T_{JO}\!\sim\!T_{JC}$ and $T_{CM}\!\sim\!T_{CO}$) may not bring significant gain. (P3) Adding too many augmentations (e.g., three or four) will be profitless or even harmful for seriously polluting the image. We thus nominate the most likely superior combinations: $T_{JOCO}$ and $T_{JCCM}$. A case of setting $T_{e}$ as $T_{A30}$ and $T_{h}$ as $T_{JOCO}$ for getting corresponding easy teacher signals $\mathbf{H}_{e}$ and hard student signals $\mathbf{H}_{h}$ is shown as below:
$\mathbf{H}_{e}=T_{A30\rightarrow A60}(f(\mathbf{I}_{e})),\quad\mathbf{I}_{e}=T_{A30}(\mathbf{I}),$ (3)
$\mathbf{H}_{h}=f(\mathbf{I}_{h}),\quad\mathbf{I}_{h}=T_{A30}(T_{CO}(T_{JO}(T_{A30}(\mathbf{I})))),~{}$
@@
where $T_{A60}$ is the default of $T_{h}$, and divided into two separate $T_{A30}$ for being compatible with $T_{e}$. To further verify the above intuitive principles, we follow the augmentations ranking way and conduct empirical studies on the performance of up to 13 selected representative combinations out of 26 choices.
As shown in Tab. III-A3 and Fig. III-A3, we can examine three principles one-by-one: (P1) $T_{MU}$ often causes adverse effects for each combination. Please refer paired combinations $c_{1}$-$c_{3}$, $c_{2}$-$c_{4}$, $c_{1}$-$c_{8}$ and $c_{2}$-$c_{9}$. Thus, we do not add it for finding superior combinations. (P2) Synergistic effects between augmentations do exist. Please refer paired combinations $c_{1}$-$c_{5}$, $c_{1}$-$c_{7}$, $c_{2}$-$c_{6}$ and $c_{2}$-$c_{7}$. Especially, the $T_{JCJO}$ with two most advanced augmentations performs the worst among combinations {$c_{1}$,$c_{2}$,$c_{5}$,$c_{6}$,$c_{7}$}, which roundly reveals the harm of violating the principle of synergy. (P3) Do not overly combine too many augmentations. Please refer paired combinations $c_{1}$-$c_{12}$, $c_{2}$-$c_{13}$, $c_{1}$-$c_{11}$ and $c_{2}$-$c_{10}$. Stacking more augmentations brings non-significant gains or even results in degradation. We attribute it to deviating from the rule of collaboration and possibly producing meaningless or difficult-to-recognize images. Based on these facts, we have sufficient reasons not to check the performance of rest combinations, and recommend two new strongest combinations $T_{JOCO}$ and $T_{JCCM}$.
TABLE I: Best mAPs of different combinations.
@@
Figure 4: The corresponding convergence curves of combinations in Tab. III-A3.
III-B Multi-path Consistency Losses
III-B1 Intuitive Motivation
$\mathcal{L}^{*}_{u}=\mathbb{E}_{\mathbf{H}_{h_{i}}\in\mathcal{H}^{n}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e}-\mathbf{H}_{h_{i}}||^{2},~{}$ (4)
where $\mathbf{H}_{e}$ and $\mathbf{H}_{h_{i}}$ are obtained as dissected in Eq. 3. The $\mathbf{H}_{e}$ keeps constant for each $\mathbf{H}_{h_{i}}$. Comparably, Data Distill [35] applies a single model to multiple transformations of unlabeled data to train a student model. Then, it ensembles predictions to obtain keypoint locations, and re-generates a pseudo heatmap for supervision. Differently, we argue that conducting a fusion on predicted heatmaps in SSHPE is harmful. We consider that there are always differences in the estimation of keypoint positions for each $\mathbf{I}_{h_{i}}$. It is an ill-posed problem to heuristically evaluate the consistency regularization contribution of each heatmap in pixel during ensemble. We will explain this in ablation studies Sec. V-C.
Despite simplicity, such a minor modification brings consistent gains over the original Single-Network under same settings. With our discovered augmentation combinations $T_{JOCO}$ and $T_{JCCM}$, the boosted Single-Network can surpass the original Dual-Network evidently. Our ablation studies reveal that the performance gain is non-trivial. We conjecture that regularizing multiple hard augmentations with a shared easy augmentation can be regarded as enforcing consistency among advanced augmentations as well, which inherits the concept of training positive-negative paired samples in contrastive learning [59, 60, 61] and SSL-related variations [62, 63, 64].
III-B2 Effectiveness Verification
As shown in Tab. III-B2 and Fig. III-B2, we also experimentally verified the major advantage of the multi-path augmentations (dubbed as MultiAugs) strategy. Here, we have two variables of MultiAugs: the number of paths and the category of augmentations. For acquisition of an optimal augmentations set, similarly, we continue with the rules summarized in the previous section, and elect 12 different MultiAugs schemes for illustrating. We can witness the effectiveness of MultiAugs from two aspects: (1) It can inherit and even expand the synergistic effects between different augmentations. (2) It can alleviate the defects caused by excessive stacking of augmentations.
TABLE II: Best mAP results of different MultiAugs tests.
@@ III-B2 Effectiveness Verification
$m_{9}$ $T_{JCCM}$ (twice) 43.6
$m_{10}$ $T_{JOCO}$ (twice) 44.0
@@
Figure 5: The corresponding convergence curves of different tests in Tab. III-B2.
@@
Specifically, by comparing $c_{1}$-$m_{1}$, $c_{2}$-$m_{2}$ and $c_{7}$-$m_{3}$, we find that MultiAugs provides comparable performance with the same augmentations. By comparing $c_{10}$-$m_{6}$, $c_{11}$-$m_{7}$, $c_{12}$-$m_{4}$ and $c_{13}$-$m_{5}$, we can observe the sustained large gains brought by MultiAugs. The scheme $m_{8}$ does not obtain a better result than $m_{11}$ showing that new augmentations $T_{JOCO}$ and $T_{JCCM}$ have their essential properties. Schemes $m_{9}$ and $m_{10}$ which utilize a single augmentation twice are also inferior to $m_{11}$ showing the cooperativity of using multi-path distinct augmentations. Moreover, the optimal scheme $m_{12}$ further unleashes capabilities of four advanced augmentations. In order to balance performance and time consumption, we do not add more augmentation paths.
III-C Analysis of Superior Augmentations
Figure 6: (a) The top-50 singular values of different augmentations. (b) The mAP vs. Entropy of different augmentations.
@@ III-C Analysis of Superior Augmentations
In this part, we tentatively analyze why employing a superior augmentation to strongly augment the unlabeled data can improve model performance. Different from UDA [36] using the improved connectivity of constructed graphs to explain, we start from the perspective of shaped feature space of unlabeled data based on the singular value spectrum, which is widely considered to be related to the model transferability and generalization [65, 66]. Specifically, we perform singular value decomposition (SVD) on features $\mathbf{F}\!\in\!\mathbb{R}^{N\times D}$ [^1] We denote $N$ as the number of samples and $D$ as feature dimensions (a.k.a, $D\leq N$). extracted by various trained models with different strong augmentations on one dataset: $\mathbf{F}\!=\!\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, where $\mathbf{U}$ and $\mathbf{V}$ is the left and right singular vector matrices, respectively, and $\mathbf{\Sigma}$ denotes the diagonal singular value matrix $\{\sigma_{1},\sigma_{2},...,\sigma_{D}\}$. Then, we plot calculated top-50 singular values in Fig. 6a. In our SVD analysis, the backbone of all six models is ResNet18. The 512-D features of 6,352 samples in COCO val-set are extracted. To further measure the flatness of the singular value distribution, we calculate the entropy of normalized singular values $\mathsf{H}_{nsv}$:
$\mathsf{H}_{nsv}=-\sum\nolimits^{D}_{i=1}\frac{\sigma_{i}}{\sum\nolimits^{D}_{j=1}\sigma_{j}}\log{\frac{\sigma_{i}}{\sum\nolimits^{D}_{j=1}\sigma_{j}}}.~{}$ (5)
@@
Usually, a larger $\mathsf{H}_{nsv}$ indicates that the feature space captures more structure in the data and thus spans more dimensions due to more discriminated representations learned. As shown in Fig. 6b, the model performance is positively correlated with the $\mathsf{H}_{nsv}$ value. Therefore, a superior augmentation facilitates better model generalization to unseen test sets.
IV Overall Framework
(a) MultiAugs with Single-Network
(b) MultiAugs with Dual-Network
@@ IV Overall Framework
Figure 7: Our proposed MultiAugs, which can utilize multiple hard augmentations and facilitate multi-path consistency training.
We exploit unlabeled images by applying multiple augmentations with integrating two key techniques introduced in Sec. III-A and Sec. III-B. Firstly, assuming that we have obtained an optimal augmentation set $\mathcal{\widehat{T}}^{n}\!=\!\{\widehat{T}_{h_{i}}|^{n}_{i=1}\}$, where $\widehat{T}_{h_{i}}$ may be an old augmentation or a novel discovered one. Then, we present our overall training framework in Fig. 7 based on either Single-Network or Dual-Network.
IV-1 MultiAugs (Single-Network)
@@
IV-2 MultiAugs (Dual-Network)
As shown in Fig. 7b, this framework learns two identical yet independent networks with each similar to the Single-Network. For one input batch in every step, each of the two networks serves as both a teacher and a student. They both are fed by easy and hard augmentations of unlabeled images $\mathbf{I}^{u}$ when they produce teacher signals and student signals, respectively. Assuming we have two networks $f_{A}$ and $f_{B}$, and also the augmented easy images $\mathbf{I}^{u}_{e}$ using $T_{e}$ and $n$-path hard images $\{\mathbf{I}^{u}_{h_{1}},\mathbf{I}^{u}_{h_{2}},...,\mathbf{I}^{u}_{h_{n}}\}$ using $\mathcal{\widehat{T}}^{n}$, we first predict the following four types of heatmaps:
$\mathbf{H}_{e,A}=T_{A30\rightarrow A60}(f_{A}(\mathbf{I}^{u}_{e})),$ (6)
$\mathcal{H}_{A}=\{\mathbf{H}_{h_{i},A}|^{n}_{i=1},\mathbf{H}_{h_{i},A}=f_{A}(\mathbf{I}^{u}_{h_{i}})\},$
$\mathbf{H}_{e,B}=T_{A30\rightarrow A60}(f_{B}(\mathbf{I}^{u}_{e})),$
$\mathcal{H}_{B}=\{\mathbf{H}_{h_{i},B}|^{n}_{i=1},\mathbf{H}_{h_{i},B}=f_{B}(\mathbf{I}^{u}_{h_{i}})\},$